# MLC at Feigenbaum points

### v2

@@ 1.1. Brief history
It took more than 10 years to prove MLC at some infinitely renormalizable parameters of bounded satellite type [DL1] (based upon the Pacmen Renormalization Theory developed in [DLS]). However, the most prominent parameter, the period-doubling Feigenbaum point corresponding to the cascade of doubling renormalizations (see Figure 1) was not covered by this result. In this paper we are filling in this gap by proving MLC at the Feigenbaum point, and in fact, in all infinitely renormalizable maps of bounded type (for which we will still preserve the same name).
@@
The universality of period-doubling cascade was discovered in the mid 70s by Feigenbaum [F1, F2] (the parameter part, see Figure 1) and independently by Coullet and Tresser [TC, CT] (the dynamical part). It is intimately related to the renormalization phenomenon in the Quantum Field Theory and Statistical Mechanics; the discovery opened up a new universality paradigm in Dynamical Systems.
@@ 1.2. Statement of the main result and its consequences
Let us consider the set ${\mathcal{I}}_{\bar{p}}$ of Feigenbaum parameters $c\in{\mathcal{M}}$ of type $\bar{p}$. Theorem A implie that each $c\in{\mathcal{I}}_{\bar{p}}$ is the intersection of the nest $M^{n}(c)$ of little $M$-copies corresponding to the $n$-fold renormalizations of $f_{c}$. Let us say that Mandelbrot set is self-similar over ${\mathcal{I}}_{\bar{p}}$ if the straightening map $\chi:{\mathcal{M}}^{n+1}(c)\rightarrow{\mathcal{M}}^{n}(\chi(c))$ (see §2.2.3) is $C^{1+{\delta}}$-conformal at any $c\in{\mathcal{I}}_{\bar{p}}$ (with uniform constants).
@@ 1.3. Rough outline of the proof
Figure 3. Primitive pull-off in the airplane combinatorics. If ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\ll{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})$, then most curves of the horizontal rectangles ${\mathcal{G}}_{1},{\mathcal{G}}_{2}$ overflow their preimages ${\mathcal{G}}^{\prime}_{1},{\mathcal{G}}^{\prime}_{2},{\mathcal{G}}^{\prime\prime}_{1}$ as illustrated. Such a configuration is impossible by the Grötzsch inequality; i.e, (1.4) holds. Peripheral preimages are omitted.
@@ 1.3.1. Pull-off in the primitive case [K, KL1]
(1.2) ${\mathcal{W}}(f_{n})={\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})+O_{p}(1)\coloneqq\sum_{\begin{subarray}{c}{\mathcal{G}}_{i}\text{ adjacent}\\ \text{to }{\mathfrak{K}}_{j}\end{subarray}}{\mathcal{W}}({\mathcal{G}}_{i})+O_{p}(1),$
i.e. the sum in (1.2) is taken over all rectangles ${\mathcal{G}}_{i}$ adjacent to ${\mathfrak{K}}_{j}$. (A “$\psi$-ql renormalization” is used to achieve (1.2), see §3.43.5.2, and (3.23).) Set also
@@
A fundamental fact is that the horizontal ${\mathcal{G}}_{i}$ are eventually (after several restrictions of the domain) aligned with the Hubbard tree $T_{f}$ of $f$ as it shown on Figure 3. Applying the Grötzsch inequality to the horizontal rectangles and their pullbacks (see the caption of Figure 3), we experience a definite loss of the horizontal weight in favor of the vertical one. The created definite vertical weight can then be pushed forward by means of the Covering Lemma [KL1]. Since all the local weights are comparable,  [^1] This observation goes back to McMullen [McM1, Theorem 9.3]. we obtain:
@@
We stress that a key combinatorial ingredient used for (1.4) is the absence of periodic horizontal rectangles; i.e. a rectangle that has an iterated lift homotopic to itself rel small Julia sets, see §2.4, §3.5.8.
@@
Figure 4. Pull-off in the Feigenbaum combinatorics for non-periodic rectangles. Either most of the horizontal degeneration is within periodic rectangles ${\mathcal{R}}_{0},{\mathcal{R}}_{1}$ such that a $(1-\delta)$-portion of each ${\mathcal{R}}_{i}$ overflows its periodic lifts ${\mathcal{R}}^{\prime}_{i}$. Or we have ${\mathcal{W}}^{\mathrm{ver}}_{\mathrm{loc}}({\mathfrak{K}}_{j})\succeq_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j}).$
@@ 1.3.2. Satellite case
Let us now assume that $f$ is an infinitely renormalizable map with the periodic-doubling combinatorics; i.e., $f$ is hybrid equivalent to a map in the period-doubling Feigenbaum combinatorial class. The general bounded case is similar.
Let ${\mathfrak{B}}\coloneqq{\mathfrak{K}}^{1}_{0}\cup{\mathfrak{K}}^{1}_{1}$ be the bouquet consisting of two level 11 11 little Julia sets of $f$. Instead of ${\mathcal{W}}(f)$, we will measure
@@
    ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\asymp_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j}){\ \ }{\ \ }{\ \ }\text{ for every }{\mathfrak{K}}_{j}$
@@
the Record Argument leads to a contradiction in case (${}^{\prime}$).
The dichotomy “Case (I) vs Case (${}^{\prime}$)” is stated as Theorem 4.1. This is a refinement of [K].
Consider now Case (I). As Figure 5 illustrates, bouquets ${\mathfrak{B}}_{0},{\mathfrak{B}}_{1}$ grow under lifting thus the “combinatorial distance” between $\widetilde{\mathfrak{B}}_{0},\widetilde{\mathfrak{B}}_{1}$ shrinks. This allows us to detect laminations ${\mathcal{S}}_{i}$ (the “differences” between ${\mathcal{R}}={\mathcal{R}}_{k}$ and its lift ${\mathcal{R}}^{\prime}$) that are much wider than ${\mathcal{W}}({\mathcal{R}})\succeq{\mathcal{W}}_{\bullet}(f_{n})$; i.e. ${\mathcal{W}}({\mathcal{S}}_{i})\geq C_{\delta}{\mathcal{W}}({\mathcal{R}})$ with $C_{\delta}\to\infty$ as $\delta\to 0$. There are two possibilities: either a substantial part of ${\mathcal{S}}_{i}$ hits preperiodic bouquets associated with $f_{n+1}$ or a substantial part of ${\mathcal{S}}_{i}$ creates a “wave” (see §5) above such a preperiodic bouquet. In the latter case, the Wave Lemma implies that ${\mathcal{W}}_{\bullet}(f_{n-1})\gg{\mathcal{W}}_{\bullet}(f_{n})$ and the Record Argument is applicable.
In the former case, we obtain ${\mathcal{W}}_{\bullet}(f_{n+1})\succeq{\mathcal{W}}({\mathcal{S}}_{i})\succeq C_{\delta}{\mathcal{W}}(f_{n})$; i.e.:
${\mathcal{W}}_{\bullet}(f_{n+1})\geq C^{\prime}_{\delta}{\mathcal{W}}(f_{n}){\ \ }{\ \ }\text{ with }{\ \ }C^{\prime}_{\delta}\to\infty\text{{\ \ }}{as}{\ \ }\delta\to 0.$
For a sufficiently small $\delta$, this eventually contradicts the Teichmüller contraction within hyperbolic classes [S1] (stated as Proposition 2.7):
@@ Acknowledgments
The first author was partially supported by the NSF grant DMS 20555322055532 20555322055532 and the ERC grant “HOLOGRAM”. The second author has been partly supported by the NSF, the Hagler and Clay Fellowships, the Institute for Theoretical Studies at ETH (Zurich), and MSRI (Berkeley).
We would like to thank Jeremy Kahn for many stimulating discussions.
@@
A number of our pictures are made with W. Jung’s program Mandel.
${\mathcal{R}}$${\mathfrak{B}}_{0}$${\mathfrak{B}}_{1}$${\mathcal{R}}^{\prime}$${\mathcal{S}}_{1}$${\mathcal{S}}_{2}$$\widetilde{\mathfrak{B}}_{0}$$\widetilde{\mathfrak{B}}_{1}$
@@ 2.1. Outline
  • ${\mathfrak{B}}^{[n]}_{i}\equiv{\mathfrak{B}}^{[n],(0)}\subset{\mathfrak{B}}^{[n],(m)}\subset{\mathfrak{K}}^{[n]}_{i}$ the associated little bushes,
@@
  • ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-1}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right)$ the union of level $n$ and height $m$ little periodic and preperiodic bushes,
@@
If ${\mathfrak{B}}_{f}={\boldsymbol{\mathfrak{K}}}^{[1]}$, i.e., ${\mathfrak{T}}_{f}\subset{\boldsymbol{\mathfrak{K}}}^{[1]}$, then we will also refer to ${\mathfrak{B}}_{f}$ as the satellite flower of $f$. Preperiodic little Julia sets and bushes are defined accordingly, 2.3.1. We will usually assume that
@@ 2.2. Quadratic-like maps
A quadratic-like map $f:X\rightarrow Y$, which will also be abbreviated as a ql map, is a holomorphic double branched covering between two Jordan disks $X\Subset Y\subset{\mathbb{C}}$. It has a single critical point that we usually put at the origin 00 0. The annulus $A=Y\setminus\overline{X}$ is called the fundamental annulus of $f\colon X\to Y$. We let $\operatorname{mod}f:=\operatorname{mod}A$. The filled Julia set ${\mathfrak{K}}_{f}$ is the set of non-escaping points:
@@
Two quadratic-like maps $f:X\rightarrow Y$ and $\tilde{f}:\tilde{X}\rightarrow\tilde{Y}$ are called hybrid conjugate if they are conjugate by a quasiconformal map $h:(Y,X)\rightarrow(\tilde{Y},\tilde{X})$ such that ¯$\bar{\partial}h=0$ a.e. on ${\mathfrak{K}}$.
@@ 2.2.2. Hubbard continua
In the superattracting case (when the crittical point is periodic) the Hubbard tree ${\mathfrak{T}}_{f}$ is defined as the smallest tree in ${\mathfrak{K}}$ containing the postcritical set such that ${\mathfrak{T}}_{f}$ intersects the components of $\operatorname{int}{\mathfrak{K}}$ along internal rays. It is invariant under $f$ and is marked with the orbit of 00 0. If ${\mathfrak{K}}$ is periodically repelling and locally connected, we can define the Hubbard dendrite ${\mathfrak{T}}_{f}$ as the smallest connected closed forward-invariant subset of ${\mathfrak{K}}$ containing the postcritical set.
@@ 2.2.5. Infinitely renormalizable maps
The ratios $q_{n}:=p_{n}/p_{n-1}$ are called relative periods. One says that such a map has bounded combinatorics of type $\bar{p}$ if the relative periods are bounded by $\bar{p}$. In this case, the map $f$ is called Feigenbaum of type $\bar{p}$. We say that a Feigenbaum map is primitive/satellite if all its renormalizations are such. A Feigenabaum map has a priori bounds if
@@ 2.3.1. Little bushes
${\boldsymbol{\mathfrak{B}}}^{[n]}\coloneqq\bigcup_{i}{\mathfrak{B}}_{i}^{[n]},{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-1}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right),{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}_{\operatorname{per}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\cap{\boldsymbol{\mathfrak{K}}}^{[n]}.$
@@
Every periodic or preperiodic little bush ${\mathfrak{B}}_{a}^{[n],(m)}$ is within a unique little periodic or preperiodic filled Julia set ${\mathfrak{K}}^{[n]}_{a}$ associated with $f_{n}$. The ${\mathfrak{B}}_{t}^{[n],(m)}$ exhaust ${\mathfrak{K}}^{[n]}_{a}$:
${\mathfrak{K}}^{[n]}_{a}=\overline{\bigcup_{m}{\mathfrak{B}}^{[n],(m)}_{a}}\ .$

### v3

@@ 1.1. Brief history
It took more than 10 years to prove MLC at some infinitely renormalizable parameters of bounded satellite type [DL1] (based upon the Pacman Renormalization Theory developed in [DLS]). However, the most prominent parameter, the period-doubling Feigenbaum point corresponding to the cascade of doubling renormalizations (see Figure 1) was not covered by this result. In this paper we are filling in this gap by proving MLC at the Feigenbaum point, and in fact, at all infinitely renormalizable parameters of bounded type (for which we will still preserve the same name).
@@
The universality of the period-doubling cascade was discovered in the mid 1970s by Feigenbaum [F1, F2] (the parameter part, see Figure 1) and independently by Coullet and Tresser [TC, CT] (the dynamical part). It is intimately related to the renormalization phenomenon in the Quantum Field Theory and Statistical Mechanics; the discovery opened up a new universality paradigm in Dynamical Systems.
@@ 1.2. Statement of the main result and its consequences
Let us consider the set ${\mathcal{I}}_{\bar{p}}$ of Feigenbaum parameters $c\in{\mathcal{M}}$ of type $\bar{p}$. Theorem B implies that each $c\in{\mathcal{I}}_{\bar{p}}$ is the intersection of the nest ${\mathcal{M}}^{n}(c)$ of little ${\mathcal{M}}$-copies corresponding to the $n$-fold renormalizations of $f_{c}$. Let us say that Mandelbrot set is self-similar over ${\mathcal{I}}_{\bar{p}}$ if the straightening map $\chi:{\mathcal{M}}^{n+1}(c)\rightarrow{\mathcal{M}}^{n}(\chi(c))$ (see §2.2.3) is $C^{1+{\delta}}$-conformal at any $c\in{\mathcal{I}}_{\bar{p}}$ (with uniform constants).
@@ 1.3. Rough outline of the proof
Figure 3. Primitive pull-off in the airplane combinatorics. If ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\ll{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})$, then most curves of the horizontal rectangles ${\mathcal{G}}_{1},{\mathcal{G}}_{2}$ overflow their preimages ${\mathcal{G}}^{\prime}_{1},{\mathcal{G}}^{\prime}_{2},{\mathcal{G}}^{\prime\prime}_{1}$ as illustrated. Such a configuration is impossible by the Grötzsch inequality; i.e., (1.4) holds. Peripheral preimages are omitted.
@@ 1.3.1. Pull-off in the primitive case [K, KL1]
(1.2) ${\mathcal{W}}(f_{n})\asymp{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})\ \coloneqq\sum_{\begin{subarray}{c}{\mathcal{G}}_{i}\text{ adjacent}\\ \text{to }{\mathfrak{K}}_{j}\end{subarray}}{\mathcal{W}}({\mathcal{G}}_{i})+O_{p}(1),$
i.e. the sum in (1.2) is taken over all rectangles ${\mathcal{G}}_{i}$ adjacent to ${\mathfrak{K}}_{j}$. (A $\psi$-ql renormalization” is used to achieve (1.2), see §3.43.5.2, and (3.23).) Set also
@@
A fundamental fact is that the horizontal ${\mathcal{G}}_{i}$ are eventually (after several restrictions of the domain) aligned with the Hubbard tree $T_{f}$ of $f$ as shown on Figure 3. Applying the Grötzsch inequality to the horizontal rectangles and their pullbacks (see the caption of Figure 3), we experience a definite loss of the horizontal weight in favor of the vertical one. The created definite vertical weight can then be pushed forward by means of the Covering Lemma [KL1]. Since all the local weights are comparable [^1] This observation goes back to McMullen [McM1, Theorem 9.3]., we obtain:
@@
We stress that a key combinatorial ingredient used for (1.4) is the absence of periodic horizontal rectangles; i.e. a rectangle that has an iterated lift homotopic to itself rel small Julia sets, see §1.3.2, §2.4, §3.5.8.
@@
Figure 4. Pull-off in the Feigenbaum combinatorics for non-periodic rectangles. Either most of the horizontal degeneration is within periodic rectangles ${\mathcal{R}}_{0},{\mathcal{R}}_{1}$ such that a $(1-\delta)$-portion of each ${\mathcal{R}}_{i}$ overflows its periodic lifts ${\mathcal{R}}^{\prime}_{i}$. Or we have ${\mathcal{W}}^{\mathrm{ver}}_{\mathrm{loc}}({\mathfrak{B}}_{j})\succeq_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{B}}_{j}).$
@@ 1.3.2. Satellite case
Let us now assume that $f$ is an infinitely renormalizable map with the period-doubling combinatorics; i.e., $f$ is hybrid equivalent to a map in the period-doubling Feigenbaum combinatorial class. The general bounded case is similar.
Let ${\mathfrak{B}}\coloneqq{\mathfrak{K}}^{1}_{0}\cup{\mathfrak{K}}^{1}_{1}$ be the bouquet consisting of two level $1$ little Julia sets of $f$. Instead of ${\mathcal{W}}(f)$, we will measure
@@
    ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{B}}_{j})\asymp_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{B}}_{j}){\ \ }{\ \ }{\ \ }\text{ for every }{\mathfrak{B}}_{j}$
@@
the Record Argument leads to a contradiction in case (II).
The dichotomy “Case (I) vs Case (II)” is stated as Theorem 4.1. This is a refinement of [K].
Consider now Case (I). As Figure 5 illustrates, bouquets ${\mathfrak{B}}_{0},{\mathfrak{B}}_{1}$ grow under lifting, thus the “combinatorial distance” between $\widetilde{\mathfrak{B}}_{0},\widetilde{\mathfrak{B}}_{1}$ shrinks. This allows us to detect laminations ${\mathcal{S}}_{i}$ (the “differences” between ${\mathcal{R}}={\mathcal{R}}_{k}$ and its lift ${\mathcal{R}}^{\prime}$) that are much wider than ${\mathcal{W}}({\mathcal{R}})\succeq{\mathcal{W}}_{\bullet}(f_{n})$; i.e. ${\mathcal{W}}({\mathcal{S}}_{i})\geq C_{\delta}{\mathcal{W}}({\mathcal{R}})$ with $C_{\delta}\to\infty$ as $\delta\to 0$. There are two possibilities: either a substantial part of ${\mathcal{S}}_{i}$ hits preperiodic bouquets associated with $f_{n+1}$ or a substantial part of ${\mathcal{S}}_{i}$ creates a “wave” (see §5) above such a preperiodic bouquet. In the latter case, the Wave Lemma implies that ${\mathcal{W}}_{\bullet}(f_{n-1})\gg{\mathcal{W}}_{\bullet}(f_{n})$ and the Record Argument is applicable.
In the former case, we obtain ${\mathcal{W}}_{\bullet}(f_{n+1})\succeq{\mathcal{W}}({\mathcal{S}}_{i})\succeq C_{\delta}{\mathcal{W}}_{\bullet}(f_{n})$; i.e.:
${\mathcal{W}}_{\bullet}(f_{n+1})\geq C^{\prime}_{\delta}{\mathcal{W}}_{\bullet}(f_{n}){\ \ }{\ \ }\text{ with }{\ \ }C^{\prime}_{\delta}\to\infty\text{{\ \ }}{as}{\ \ }\delta\to 0.$
For a sufficiently small $\delta$, this eventually contradicts the Teichmüller contraction within hybrid classes [S1] (stated as Proposition 2.7):
@@ Acknowledgments
The first author was partially supported by the NSF grant DMS $2055532$ and the ERC grant “HOLOGRAM”. The second author has been partly supported by the NSF, the Hagler and Clay Fellowships, the Institute for Theoretical Studies at ETH (Zurich), and MSRI (Berkeley).
We would like to thank Jeremy Kahn for many stimulating discussions. We are grateful to the referee for a comprehensive reading of the manuscript. A number of our pictures are made with W. Jung’s program Mandel.
@@
[…]
@@ 2.1. Outline
  • ${\mathfrak{B}}^{[n]}_{i}\equiv{\mathfrak{B}}^{[n],(0)}_{i}\subset{\mathfrak{B}}^{[n],(m)}_{i}\subset{\mathfrak{K}}^{[n]}_{i}$ the associated little bushes,
@@
  • ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right)$ the union of level $n$ and height $m$ little periodic and preperiodic bushes,
@@
If ${\mathfrak{B}}_{f}={\boldsymbol{\mathfrak{K}}}^{[1]}$, i.e., ${\mathfrak{T}}_{f}\subset{\boldsymbol{\mathfrak{K}}}^{[1]}$, then we will also refer to ${\mathfrak{B}}_{f}$ as the satellite flower of $f$. Preperiodic little Julia sets and bushes are defined accordingly, §2.3.1. We will usually assume that
@@
Periodic little Julia sets will be usually labeled dynamically:
$f\left({\mathfrak{K}}^{[n]}_{i}\right)={\mathfrak{K}}^{[n]}_{i+1},{\ \ }{\ \ }{\ \ }i,\ i+1\in{\mathbb{Z}}/{p_{n}}\simeq\{0,1,\dots,p_{n}-1\}.$
However, in the satellite combinatorics, we will often label the ${\mathfrak{K}}^{[n]}_{i}$ with respect to their cyclic order in the satellite flower; see, for instance, Lemma 2.3. Preperiodic little Julia sets ${\mathfrak{K}}^{[n]}_{j}$ will be labeled with some subindex. The labeling of little bushes is naturally inherited from the labeling of little Julia sets: ${\mathfrak{B}}^{[n],(m)}_{i}\subset{\mathfrak{K}}^{[n]}_{i}$. We will always indicate when the cyclic order is used. And we will also mention when a preperiodic object appears.
@@ 2.2. Quadratic-like maps
A quadratic-like map $f:X\rightarrow Y$, which will also be abbreviated as a ql map, is a holomorphic double branched covering between two Jordan disks $X\Subset Y\subset{\mathbb{C}}$. It has a single critical point that we usually put at the origin $0$. The annulus $A=Y\setminus\overline{X}$ is called the fundamental annulus of $f\colon X\to Y$. We let $\operatorname{mod}f:=\operatorname{mod}A$. The filled Julia set ${\mathfrak{K}}_{f}$ is the set of non-escaping points:
@@
Two quadratic-like maps $f:X\rightarrow Y$ and $\tilde{f}:\tilde{X}\rightarrow\tilde{Y}$ are called hybrid conjugate if they are conjugate by a quasiconformal map $(Y~,X~)\displaystyle h$$\rightarrow(\tilde{Y},\tilde{X})$ such that $\bar{\partial}h=0$ a.e. on ${\mathfrak{K}}$.
@@ 2.2.2. Hubbard continua
In the superattracting case (when the critical point is periodic) the Hubbard tree ${\mathfrak{T}}_{f}$ is defined as the smallest tree in ${\mathfrak{K}}$ containing the postcritical set such that ${\mathfrak{T}}_{f}$ intersects the components of $\operatorname{int}{\mathfrak{K}}$ along internal rays. It is invariant under $f$ and is marked with the orbit of $0$. If ${\mathfrak{K}}$ is periodically repelling and locally connected, we can define the Hubbard dendrite ${\mathfrak{T}}_{f}$ as the smallest connected closed forward-invariant subset of ${\mathfrak{K}}$ containing the postcritical set.
@@ 2.2.5. Infinitely renormalizable maps
The ratios $q_{n}:=p_{n}/p_{n-1}$ are called relative periods. One says that such a map has bounded combinatorics of type $\bar{p}$ if the relative periods are bounded by $\bar{p}$. In this case, the map $f$ is called Feigenbaum of type $\bar{p}$. We say that a Feigenbaum map is primitive/satellite if all its renormalizations are such. A Feigenbaum map has a priori bounds if
@@ 2.3.1. Little bushes
${\boldsymbol{\mathfrak{B}}}^{[n]}\coloneqq\bigcup_{i}{\mathfrak{B}}_{i}^{[n]},{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right),{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}_{\operatorname{per}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\cap{\boldsymbol{\mathfrak{K}}}^{[n]}.$
@@
Every periodic or preperiodic little bush ${\mathfrak{B}}_{a}^{[n],(m)}$ is within a unique little periodic or preperiodic filled Julia set ${\mathfrak{K}}^{[n]}_{a}$ associated with $f_{n}$. The ${\mathfrak{B}}_{a}^{[n],(m)}$ exhaust ${\mathfrak{K}}^{[n]}_{a}$:
${\mathfrak{K}}^{[n]}_{a}=\overline{\bigcup_{m}{\mathfrak{B}}^{[n],(m)}_{a}}\$.